OpenStudy (anonymous):

3 + 2x = -3 3 - 2x what does x equal?

OpenStudy (anonymous):

What does 3-2x equal to in your question?

OpenStudy (anonymous):

\[\frac{3+2x}{3-2x}\]?

OpenStudy (anonymous):

finish your question please

OpenStudy (anonymous):

What is this i don't even.

OpenStudy (anonymous):

is it how Thomas has written it?

OpenStudy (anonymous):

Assuming Thomas's way: 3 + 2x = -3(3-2x) 3 + 2x = -9 + 6x 12 = 4x x = 3

OpenStudy (anonymous):

Since the answer is clean when interpreted as above, I would imagine that is the correct form.